Suppose you invested $50,000 into an account at a private bank that attracts annual

interest of 20% using this formula

A= Pe^rt . How long would it take to ensure it

grows to become $88,000?


1
Expert's answer
2021-06-17T12:47:47-0400

"A=P×e^{r×t}\\88000=50000×e^{0.2×t}\\\frac{80000}{50000}=2.718282^{0.20×t}\\0.2×t\space log(2.718282)=log\frac{44}{25}\\0.20×t=\frac{log \frac{44}{25}}{log(2.718282)}\\t=2.82657"